OpenStudy (anonymous):

A swimming pool can be filled in 18 hours if water enters through a pipe alone, or in 23 hours if water enters through a hose alone. If water is entering through both the pipe and the hose, how long will it take to fill the pool?

OpenStudy (anonymous):

Let it take t hours,then \[\frac{ 1 }{ t }=\frac{ 1 }{ 18 }+\frac{ 1 }{23 },find t\]